OpenStudy (anonymous):

Solve the following system of equations: 10p + 9q = 3(2p + q) + 24 4/3p + 2q - 8 =0 a) (6,6) b) (-6,-6) c) always true d) none of these When I did this, I got q=0 and p=6

OpenStudy (anonymous):

Are the following the equations?\[\left\{10 p+9 q=3 (2 p+q)+24,\frac{4 p}{3}+2 q-8=0\right\} \]

OpenStudy (anonymous):

no. the fractions are there, and then the variable. (4/3)p